Your company pays $30,000 for the 5,000 megawatts of electricity it used this month. To make...

  1. Your company pays $30,000 for the 5,000 megawatts of electricity it used this month. To make your product, your company performs Activity A, Activity B, and Activity C. If Activity A used up 1,250 megawatts of electricity, Activity B used up 2,750 megawatts of electricity, and Activity C used up 1,000 megawatts of electricity, then how much of the electricity bill should be assigned to
    1. Activity A cost pool?
    2. Activity B cost pool?
    3. Activity C cost pool?

  1. Your company makes Product 1 and Product 2 using Activity A and Activity B. The cost driver for Activity A is direct labor hours and the cost driver for Activity B is number of setups. Activity A costs $200,000 and Activity B costs $600,000. If Product 1 requires 800 direct labor hours and 70 setups, but Product 2 requires 400 direct labor hours and 90 setups, then what is
    1. the activity rate for Activity A?
    2. the activity rate for Activity B?
    3. Product 1’s activity proportion for Activity A?
    4. Product 2’s activity proportion for Activity A?
    5. Product 1’s activity proportion for Activity B?
    6. Product 2’s activity proportion for Activity B?
Answer #1

1)

Cost pool = Total cost * Activity consumed / Total activity

Activity A cost pool

= 300,000*1,250/5,000

=$75,000

Activity B cost pool

= 300,000*2,750/5,000

=$165,000

Activity C cost pool

= 300,000*1,000/5,000

= $60,000

2)

Activity A cost = $200,000

Activity B Cost = $600,000

Activity A uses Direct labour hours

Activity B uses Number of set ups

Product 1 uses 800 direct labour hours and 70 setups

Product 2 uses 400 direct labour hours and 90 setups

A)

Activity rate for activity A

= Total cost / Total direct labour hours

= 200,000/(800 + 400)

= $166.67

B)

Activity rate for activity B

= Total cost / Total number of setups

= 600,000(90 + 70)

= $3,750

C)

Product 1 proportion for Activity A

= 800*166.67

= $133,333

D)

Product 2 proportion for Activity A

= 400*166.67

=$66,667

E)

Product 1 proportion for Activity B

= 70*3,750

=$262,500

Product 2 proportion for Activity B

= 90*3,750

= $337,500

If you find the answer helpful please upvote